suppose A and B are independent events. If P(A|B)=0.25 and P(B)=0.4, what is P(A U B)?
Answer by Edwin McCravy(20055)

suppose A and B are independent events. If P(A|B)=0.25 and P(B)=0.4,
what is P(A U B)?
Since A and B are independent, P(A|B) = P(A) = 0.25, P(A&B)=P(A)P(B)

P(A U B) = P(A) + P(B) - P(A&B)

P(A U B) = P(A) + P(B) - P(A)P(B)

P(A U B) = 0.25 + 0.4 - (0.25)(0.4)

P(A U B) = 0.25 + 0.4 - 0.1

P(A U B) = 0.55

Edwin